OpenStudy (anonymous):

Solve for x. 5x - 11 = 7x - 3 x = -7 x = -4 x = 2 ---- 3 x = 7 --- 3

OpenStudy (anonymous):

5x - 11 = 7x - 3 -7x -7x

OpenStudy (anonymous):

-2x-11=-3

OpenStudy (anonymous):

-2x-11=-3 +11 +11

OpenStudy (purplemexican):

-4

OpenStudy (anonymous):

-.- I was trying to explain first....

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

-2x=8

OpenStudy (anonymous):

divide both sides by -2

OpenStudy (anonymous):

to get x=-4

OpenStudy (anonymous):

purplemexican was correct

OpenStudy (anonymous):

but he/she didn't show the steps

OpenStudy (purplemexican):

i was gonna put a link to wolframalpha but im to tired sorry